OpenStudy (anonymous):

Solve the 2nd order non-homogenous ODE Initial value problem: y''(t)+y(t) = sec(t) y(0)=1 , y'(0)=2

OpenStudy (anonymous):

Anyone familiar with Diff Eq?

OpenStudy (jamesj):

This is slightly tricky question. Have your learnt yet about turning 2nd order equations into a system of first order equations? If not, what other methods have you learnt?

OpenStudy (anonymous):

im up to Laplace transform however this is in the chapter review for undefined coefficients.

OpenStudy (jamesj):

Ok. So use that method. The homogeneous solutions of course are sin x and cos x, hence the particular solution is of the form yp = u(x) sin x + v(x) cos x. Now grind away ...

OpenStudy (anonymous):

yeah i got the homogeneous sol'n however i dont know where to start after that... nothing in this chapter deals with sec,csc,tan,cot.... so do i just treat it as 1/cos?

OpenStudy (anonymous):

the homogenous sol'n i got was: \[c _{1}e ^{0t}\cos(-t)+c _{2}e ^{0t}\sin(-t)\]

OpenStudy (amistre64):

y''+y = 0 ; roots are complex; r(1+i1, 1-i1) \[y=e^t(c_1cos(t)+c_2sin(t))\] right?

OpenStudy (jamesj):

No. y = c1.cos t + c2.sin t

OpenStudy (amistre64):

\[r(a+ib,a-ib)\] \[y=e^{ax}(c_1cos(bx)+c_2sin(bx))\] is what i recall

OpenStudy (anonymous):

amistre those roots are incorrect. using quad. equations since 'b' = 0 there is no real part asociated

OpenStudy (amistre64):

hmmm ..... im open to being wrong :)

OpenStudy (jamesj):

Amistre, for heaven's sake, with y = e^(rt), the characteristic equation is r^2 + 1 = 0 and therefore r = +-i.

OpenStudy (amistre64):

oh ;)

OpenStudy (anonymous):

but the main problem is i dont know what to 'guess' as the y_p

OpenStudy (amistre64):

\[r^2+1=(r+i)(r-i)\] i had the right thought, my fingers just betrayed me lol

OpenStudy (anonymous):

would it be Asec(t)+Bcsc(t)?

OpenStudy (asnaseer):

shouldn't the characteristic equation be r^2+r=0 which gives r=0 or -1?

OpenStudy (jamesj):

No.

OpenStudy (amistre64):

r^2 = y'' r = y' c = y

OpenStudy (amistre64):

that gets me at times too

OpenStudy (asnaseer):

sorry mis-read the equation i thought the 2nd term had a prime on it - must change my glasses! :-)

OpenStudy (jamesj):

Now, the guess. I think you mean btw, the method of undetermined coefficients. Try yp = A sec x + B sec x.tan x + C cos x + D cos x.ln(cos x)

OpenStudy (amistre64):

could try sin(t) for the other part .... maybe. just did a few of these before bedtime last night, but they werent triged up

OpenStudy (jamesj):

Actually, just yp = A sec x + B sec x.tan x + C cos x.ln(cos x)

OpenStudy (anonymous):

where do you get those from? just need to learn it or?

OpenStudy (anonymous):

because it is no where to be found in my book

OpenStudy (jamesj):

sec x comes from the in homogenous part sec x . tan x from its derivative Now more speculative, what happens when you try and integrate by parts tan x * something, you get something * -ln(cos x) So sec x.ln(cos x) and cos x.ln(cos x) makes sense

OpenStudy (jamesj):

So actually, last iteration: try yp = A sec x + B sec x.tan x + C cos x.ln(cos x) + D sec x.ln(cos x)

OpenStudy (amistre64):

yp = A cos(t) + B sin(t) yp' = A' cos(t) - Asin(t) + B'sin(t) +Bcos(t) assume: A'cos(t) + B'sin(t) = 0 cant recall y yp' = - Asin(t) +Bcos(t) yp'' = -A'sin(t)-Acos(t) +B'cos(t) -Bsin(t) yp''+y = -A'sin(t)-Acos(t) +B'cos(t) -Bsin(t) +A cos(t) + B sin(t) yp''+y = -A'sin(t)+B'cos(t) = csc(t)

OpenStudy (amistre64):

then solve the system of equations from: A'cos(t) + B'sin(t) = 0 -A'sin(t)+B'cos(t) = csc(t)

OpenStudy (amistre64):

B' = csc.cos/ cos^2+sin^2 = cot

OpenStudy (amistre64):

A' = 1/-1 = -1

OpenStudy (amistre64):

i hope i did those crosses right

OpenStudy (jamesj):

Ok: so what amistre's written is what I thought you meant in terms of method when you said "undefined coefficients", which goes under the usual name "Variation of parameters". But you are using "Method of Undetermined Coefficients" which is why you need a good guess for the particular solution.

OpenStudy (anonymous):

yeah been doing alot of this my brain is starting to go haywire lol

OpenStudy (amistre64):

int(A')dt = int(-1)dt A = -t int(B')dt = int(cot(t))dt B = ln|sin(t)|

OpenStudy (jamesj):

So which one do you mean? Variation of Parameters or Undetermined Coefficients?

OpenStudy (anonymous):

undetermined coefficents

OpenStudy (anonymous):

amistre i dont see how you went from yp = A cos(t) + B sin(t) to that yp' = A' cos(t) - Asin(t) + B'sin(t) +Bcos(t) shouldnt y_p'= -aCos(t)+bSin(t)?

OpenStudy (amistre64):

y = yh + yp \[y=c_1cos(t)+c_2sin(t)-t\ cos(t)+ln|sin(t)|\ sin(t)\] ill rechk it, but in the end I get this; unless im forgeting another c

OpenStudy (amistre64):

yp = A cos(t) + B sin(t) yp' = A' cos(t)+A cos(t)' +B' sin(t) + B sin(t)' yp' = A' cos(t)-A sin(t) +B' sin(t) + B cos(t)

OpenStudy (anonymous):

but A and B are constants. why would you use a' and b' ?

OpenStudy (amistre64):

they arent constants they are unknown functions

OpenStudy (amistre64):

we are trying to find the unknown functions A(t) and B(t)

OpenStudy (jamesj):

amistre: you're not using the method he wants to use.

OpenStudy (amistre64):

im aware of that ;) but its still good practice

OpenStudy (anonymous):

well that explains why i couldn't follow along lol

OpenStudy (anonymous):

what is the name of that method?

OpenStudy (amistre64):

even if we use tham as variables, they are still unknowns that get 'ed

OpenStudy (amistre64):

cant recall

OpenStudy (jamesj):

The method amis is using is called Variation of Parameters. I'll tell you flat out that Undetermined Coeffs is not a great method here as it's not obvious what form of the answer to use.

OpenStudy (jamesj):

And I'll also tell you the answer so you can work towards it: yp = cos x . ln(cos x)

OpenStudy (anonymous):

ok. that makes so much more sense (looking at the next section)

OpenStudy (anonymous):

midterm in 24hrs... gotta keep studying lol.

OpenStudy (jamesj):

err, for the record, who is on the poster you're holding up in your photo?

OpenStudy (anonymous):

osama bin ladin haha was at gun range in vegas

OpenStudy (jamesj):

got it.

OpenStudy (anonymous):

thanks for the help

OpenStudy (amistre64):

i apparently cant factor r^2+1 correctly, but I can recall pretty much how to do variation param.... you know how much algebra i had to forget to learn that?